# RealmDreamer: Text-Driven 3D Scene Generation with Inpainting and Depth Diffusion

### v1

@@ A.1 Additional Implementation Details
We intend to open-source our code later on. In addition, we describe some key implementation details to assist reproducibility.
A.1.1 Point Cloud Generation
@@
A.1.2 Occlusion Volume Computation
@@
Figure 10: We ablate the importance of DDIM Inversion. Like other methods [31], we find that the inversion allows more details to be synthesized by our method.
A.1.3 Optimization
Use of DDIM Inversion. During the inpainting and finetuning stage (Sec 4.2, 4.3 in the original paper), we find it helpful to obtain the noisy latent $z_{t}$ using DDIM inversion [59], where $z=\mathcal{E}(x)$, $x$ is the rendered image, and $t$ is a timestep corresponding to the amount of noise added. This is similar to prior work on 2D/3D editing and synthesis using pre-trained diffusion models [20, 31]. We demonstrate the importance of doing so in Fig. 10, where DDIM inversion can significantly improve the detail in the optimized model. During the inpainting stage, we use 25 steps to sample an image from pure noise, and during finetuning, we use 100 steps.
Hyperparameter Weights. We set $\lambda_{\text{latent}}=0.1,\lambda_{\text{anchor}}=10000$ during the inpainting stage, and $\lambda_{\text{latent}}=0.01,\lambda_{\text{anchor}}=0$ during the finetuning stage. The other parameters are set as $\lambda_{\text{image}}=0.01,\lambda_{\text{lpips}}=100,\lambda_{\text{depth}}=1000\text{, and }\lambda_{\text{opacity}}=10$.
@@ A.1.3 Optimization
Gaussian Splatting. We initialize our gaussian splatting model during the inpainting stage, using the point cloud from stage 1, where each point is an isotropic gaussian, with the scale set based on the distance to its nearest neighbors. During the inpainting stage, we use a constant learning rate of $0.01$ for rotation, $0.001$ for the color, and $0.01$ for opacity. The learning rate of the geometry follows an exponentially decaying scheduler, which decays to $0.00005$ from $0.01$ over $100000$ steps, after $5000$ warmup steps. Similarly, the scale is decayed to 0.0001 from $0.005$ over $10000$ steps, after $7000$ warmup steps. During the finetuning stage, we use a constant learning rate of $0.01$ for rotation, $0.001$ for the color, $0.01$ opacity, and $0.0001$ for scale. We use an exponentially decaying scheduler for the geometry, which decays to $0.0000005$ from $0.0001$ over $3000$ steps, after $750$ warmup steps.
A.1.4 Comparison with baselines.
Text2Room [23] We use the official implementation of Text2Room on Github. To ensure a fair comparison, we estimate depth using Marigold [29] and DepthAnything [69] as in our technique, replacing the original IronDepth [2]. The rest of the pipeline is kept the same.
A.2 Additional Qualitative Results

### v2

@@ Appendix D Additional Implementation Details
We intend to open-source our code upon publication. In addition, we describe some key implementation details to assist reproducibility.
D.1 Point Cloud Generation
@@
D.2 Occlusion Volume Computation
@@
D.3 Optimization
Hyperparameter Weights. We set $\lambda_{\text{latent}}=0.1,\lambda_{\text{anchor}}=10000$ during the inpainting stage, and $\lambda_{\text{latent}}=0.01,\lambda_{\text{anchor}}=0$ during the refinement stage. The other parameters are set as $\lambda_{\text{image}}=0.01,\lambda_{\text{lpips}}=100,\lambda_{\text{depth}}=1000\text{, and }\lambda_{\text{opacity}}=10$.
@@ D.3 Optimization
Opacity Loss. We compute the opacity loss as the binary cross entropy of each splat’s opacity $\sigma_{i}$ with itself. This encourages the opacity to reach either 0 or 1.
@@
Gaussian Splatting. We initialize our gaussian splatting model during the inpainting stage, using the point cloud from stage 1, where each point is an isotropic gaussian, with the scale set based on the distance to its nearest neighbors. During the inpainting stage, we use a constant learning rate of $0.01$ for rotation, $0.001$ for the color, and $0.01$ for opacity. The learning rate of the geometry follows an exponentially decaying scheduler, which decays to $0.00005$ from $0.01$ over $100000$ steps, after $5000$ warmup steps. Similarly, the scale is decayed to 0.0001 from $0.005$ over $10000$ steps, after $7000$ warmup steps. During the refinement stage, we use a constant learning rate of $0.01$ for rotation, $0.001$ for the color, $0.01$ opacity, and $0.0001$ for scale. We use an exponentially decaying scheduler for the geometry, which decays to $0.0000005$ from $0.0001$ over $3000$ steps, after $750$ warmup steps. During the inpainting distillation, we also dilate $M_{\text{occl}}$ to improve cohesion at mask boundaries. Further, we find it essential to mask the latent-space L2 loss, to prevent unwanted gradients outside the masked region.
Figure 11: Comparison of sampling from 2D inpainting models and our optimized model. Left: Renders from the point cloud generated in stage 1. Middle (cols 2-4): Inpainted Samples of the previous render using an occlusion-based inpainting mask and Stable Diffusion [50]. Right: A render from our final 3DGS model for the corresponding scene. We find that our distillation techniques produce results with high cohesion while avoiding many artifacts from ancestral sampling of 2D inpainting models.
Appendix E User Study
For comparison with ProlificDreamer [64], DreamFusion [44], and LucidDreamer [14], we showed participants side-by-side videos comparing our method to the baseline. For fairness, we use the same camera trajectory in all videos. The order of the videos was also randomized to prevent any biases due to the order of presentation. The user’s preference was logged along with a brief explanation.
When comparing with Text2Room [26], instead of a video, we showed users side-by-side sets of three multiview images for each prompt, due to the degeneracy of the output mesh far from the starting camera pose. The user’s preferred triplet was logged along with their brief explanation. The images we showed looked slightly left and right of the reference pose $P_{\text{ref}}$.
E.1 Common themes of the user study.
ProlificDreamer [64] can produce cloudy results. Several participants described the NeRF renders as containing “moving clouds”, a “hazy atmosphere”, and a “blotch of colours”. This can likely be attributed to the presence of floaters in the model, which is evident in the noisy depth maps shown in Appendix H. In contrast, participants described our method as “clean and crisp when it comes to the colors and sharpness of the pixels” and looking realistic, without the presence of over-saturated colors.
LucidDreamer [14]’s scenes lack cohesion and can be distorted. Multiple participants pointed out that LucidDreamer’s scenes degrade in quality when moving away from the initial pose. One participant wrote “The image on the left loses cohesion when rotated.” referring to LucidDreamer and in contrast another wrote “There is less visual distortion when the camera is moved around the room.” about RealmDreamer. Some participants also noted that objects produced by our technique were more solid, with one participant noting “The shapes are solid on the right and hold their form.”. These comments underscore the limitations of purely iterative approaches.
Appendix F Additional Discussion
F.1 Impact of Distillation
In Fig. 11, we show the importance of our distillation process for filling in occluded regions and the challenge in doing so. Column 1 shows renders following stage 1, which contains large holes, giving objects a thin look (such as the bear in row 2 or the table in row 1). By computing an occlusion volume and obtaining inpainting masks, we can inpaint these renders to obtain several inpainted samples (columns 2-4). However, these samples can contain several artifacts. For instance, in row 1 of Fig. 11, the surface of the table is quite cluttered in individual samples. This is likely due to the challenge of inpainting images with complex masks that are out of distribution. These images also show the challenge in building cohesive scenes with single view inpainting. For instance the blackboard in row 2 has multiple shades of green in the 2D samples. Despite these challenges, our final render for the scene, in column 5 of Fig. 11 is clean and free of stray artifacts such as bright colours or ambiguous objects. We attribute this difference to our distillation process.
As mentioned earlier, since we optimize over multiple views, we are less susceptible to artifacts present in individual samples and can produce 3D inpaintings that satisfy multiple views. Prior work, such as Text2Room [26] instead relies primarily on dilating masks and deleting regions of generated scenes to simplify the inpainting process. Our inpainting distillation process does not require any aggressive modification to the scene but can produce high-quality results. We highly encourage the viewer to view the video renderings to appreciate the extent of occluded regions that our distillation technique generates.
Appendix G Limitations
Figure 12: Janus Problem due to multi-view optimization. Since we optimize over multiple-views, sometimes the final model can show the same object multiple times to satisfy all views, such as the pair of glasses above the octopus. Prompt: “A blue octopus wearing glasses on a couch in the living room, watercolor style”
Janus Problem. By adopting a distillation based approach, we occasionally encounter the Janus problem, where the face of an object appears multiple times across renders. An example is shown in Fig. 12 As we focus on scene generation and additionally condition on the 3D scene, this is less pronounced than in object generation [44] and can likely be alleviated with view-dependent prompting [2].
Artifacts in rendering. Some scenes also display artifacts at the surface of objects over a wide baseline. We believe improvements to our 3DGS implementation, such as by incorporating anti-aliasing, and surface regularizers might help with this. We note that our results are still significantly better than prior work and uses only 2D priors.
Appendix H Additional Qualitative Results
@@
Figure 13: Prompt: ”An astronaut in a cave, trending on artstation, 8k image”
Figure 14: Prompt: ”Editorial Style Photo, Coastal Bathroom, Clawfoot Tub, Seashell, Wicker, Mosaic Tile, Blue and White”
Figure 15: Prompt: ”A minimalist bedroom, 4K image, high resolution”
Figure 16: Prompt: ”A boy sitting in a boat in the middle of the ocean, under the milkyway, anime style”
Figure 17: Prompt: ”a living room, high quality, 8K image, photorealistic”
Figure 18: Prompt: ”A marble bust in a museum with pale teal walls, framed paintings, marble patterned floor, 4k image”
Figure 19: Prompt: ”A bear sitting in a classroom with a hat on, realistic, 4k image, high detail”
Figure 20: Prompt: ”An old car overgrown by vines and weeds, high quality image, photorealistic, 4k image”
Figure 21: Prompt: ”Small lavender room, soft lighting, unreal engine render, voxels.”
Figure 22: Prompt: ”White grand piano on wooden floors in an empty hall, 4k image”
Figure 23: Prompt: ”A highly detailed image of the resolute desk in the oval office, 4k image”
Figure 24: Prompt: ”A bohemian living room, colorful textiles, vibrant, eclectic, 4k image, photorealistic”
Figure 25: Prompt: ”Retro arcade room with posters on the walls, retro art style, illustration”
Figure 26: Prompt: ”A thick elven forest, fantasy art, landscape, picturesque, 4k image”
Figure 27: Prompt: ”A sunny royal traditional Japanese bedroom, 4k image, ornate, high detail”
Figure 28: Prompt: ”An old charming stone kitchen, 4k image, photorealistic, high detail”
Figure 29: Prompt: ”Fantasy lighthouse in the Arctic, surrounded by a world of ice and snow, shining with a mystical light under the aurora borealis, 4k, sharp”
Figure 30: Prompt: ”A steampunk bedroom with glass ceilings, photorealistic, 4k image, bright lighting”
Figure 31: Prompt: ”A majestic peacock, surfing a tall wave, photorealistic, detailed image, 4k image”
Figure 32: Prompt: ”A victorian living room with a grand fireplace and a long sofa, painting over the fireplace, mysterious vibe, giant windows, 4k image, photorealistic”